Calculate the bond issue price for the following scenarios and state whether it would be issued...

  1. Calculate the bond issue price for the following scenarios and state whether it would be issued at a discount, a premium or at face value. (The interest rates showing are the stated interest rates.) (2pts each):

a. $300,000 8% bond issued at 100

b. $900,000 4% bond issued at 105

c. $600,000 5% bond issued at 93
Answer #1

Issue Price = Face Value * Issue Rate

Discount or Premium = Bond Price -Issue Price

Scenario Face Value Issued AT Issue Price (Answers ) Premium or Discount ( Answers )
A 300,000 100 = 300,000*1 =        300,000 AT Face Value
B 900,000 105 =900,000*1.05 =   $945,000 Premium of 45,000
C 600,000 93 = 600,000*0.93=   $558,000 Discount of $ 42,000